# A Comprehensive Evaluation of Large Language Models on Mental Illnesses

### v1

@@ 2.4 Models
Our model selection encompassed both closed-source models, accessible through APIs, and open-source models, which offer greater transparency and flexibility for research purposes. We included models from various families, including OpenAI’s GPT series, Anthropic’s Claude models, Mistral AI’s models, Google’s models, Meta’s Llama series, and finally, Phi-2 and MentaLlama, a model specifically fine-tuned on mental health-related data. Additionally, we attempted to include Gemini 1.5 Pro in our evaluation however due to rate limit constraints, and challenges mentioned in 5, we were not able to include it.
Table 2 provides a summary of the models used in our experiments, along with their respective sizes and the API s through which we accessed them.
Table 2: Summary of large language models evaluated in this analysis
Model Size (Parameters) Source API Service Provider
Claude 3 Haiku [52] Proprietary Anthropic Claude API
Claude 3 Sonnet [52] Proprietary Anthropic Claude API
Claude 3 Opus [52] Proprietary Anthropic Claude API
Claude 3.5 Sonnet [53] Proprietary Anthropic Claude API
Gemma 2b [54] 2B Google Nvidia NIM
Gemma 7b [54] 7B Google Groq
Gemma 2 9b [55] 9B Google Nvidia NIM
Gemma 2 27b [55] 27B Google Nvidia NIM
Gemini 1.0 Pro [8] Proprietary Google Gemini API
GPT-4o mini [56] Proprietary OpenAI OpenAI API
GPT-3.5-Turbo [57] 175B OpenAI OpenAI API
GPT-4o [58] Proprietary OpenAI OpenAI API
GPT-4-Turbo [7] Proprietary OpenAI OpenAI API
GPT-4 [7] Proprietary OpenAI OpenAI API
Llama 3 8b [59] 8B Meta Groq
Llama 3.1 8b [59] 8B Meta Together AI
Llama 2 13b [60] 13B Meta Together AI
Llama 2 70b [60] 70B Meta Together AI
Llama 3 70b [59] 70B Meta Groq
Llama 3.1 70b [59] 70B Meta Together AI
Llama 3.1 405b [59] 70B Meta Together AI
Mistral 7b [61] 7B Mistral AI Together AI
Mistral NeMo [62] 12B Mistral AI Mistral AI API
Mixtral 8x7b [63] 45B (13B) Mistral AI Groq
Mixtral 8x22b [63] 141B (39B) Mistral AI Together AI
Mistral Medium [64] Proprietary Mistral AI Mistral AI API
Mistral Large [65] Proprietary Mistral AI Mistral AI API
Mistral Large 2 [66] Proprietary Mistral AI Mistral AI API
Phi-2 [67] 2.7b Phi-1 Labs Nvidia NIM
Phi-3-mini [68] 3.8B Phi-1 Labs Nvidia NIM
Phi-3-small [68] 7B Phi-1 Labs Nvidia NIM
Phi-3-medium [68] 14B Phi-1 Labs Nvidia NIM
Qwen 2 72b [69] 72B Alibaba Cloud Together AI
Notes: The Mixtral 8x7b and 8x22b models have two sizes because the total size and the number of parameters utilized at any point in inference are different.
@@ 2.5 Prompt Templates
The prompt, or instruction, given to an LLM plays a great role in its performance on a given task. Prior research has shown that even slight variations in prompt wording can significantly impact the quality of LLM outputs [70]. Therefore, we dedicated considerable effort to crafting effective prompts tailored to each task in our study.
To optimize LLM performance on social media post analysis, we employed prompts that encouraged the models to adopt the role of a psychiatrist. Role-playing has proven effective in eliciting relevant and contextually appropriate responses from LLM s [71, 72]. We also utilized the concept of prompt templates, which consist of a fixed structure with variable elements that can be adapted for different tasks. This approach ensured consistency across experiments while minimizing prompt-based bias.
@@ 2.5.1 Binary Disorder Detection Prompt Templates
As shown in 3, Prompts BIN-1 and BIN-2 utilized a straightforward structure, first instructing the LLM to assume the role of a psychiatrist and then directly asking whether the poster of a given social media post exhibited a specific disorder. We included a final instruction to elicit a "yes" or "no" response for easier parsing of the LLM output.
@@
[Task] Act as a psychologist. Label the writer of the provided social media post as ’depressed’ or ’not depressed’ according to provided guidelines. [Guidelines] - Concise Response: Respond only with ’Yes’ (depressed) or ’No’ (not depressed). - No Explanations: Don’t provide explanations for your assessment. - Ambiguity: If the post is unclear, choose the most probable label. [Post]
@@
Prompts BIN-3 and BIN-4 introduced a more structured format, explicitly separating the prompt into "Task," "Guidelines," and "Post" sections. The guidelines emphasized the desired output format (yes/no), discouraged explanations, and instructed the LLM to choose the most probable label in cases of ambiguity. This structure aimed to improve the LLM s’ adherence to the task instructions and reduce the incidence of irrelevant or unparseable responses. The inspiration for this structure was from Gemini, as this is the structure used for its training [8].
@@ 2.5.2 Disorder Severity Evaluation Prompt Templates
Figure 1: Severity Task Prompts
@@ 2.6 Evaluation Metrics
For the disorder severity evaluation task, we employed two metrics to assess the performance of LLM s: BA and Mean Absolute Error (MAE). BA, in this context, refers to the proportion of posts for which the LLM predicted the exact severity level assigned by human annotators. It serves as a measure of how often the model gets the severity classification exactly right. On the other hand, MAE quantifies the average magnitude of the errors in the predicted severity ratings. It provides a complementary perspective on model performance by assessing how close, on average, the predicted severity levels are to the true levels, even if they are not exactly correct. A model could potentially improve in one metric while worsening in the other.
@@ 2.7 Parsing Model Outputs
We observed that model adherence to instructions varied widely and did not necessarily correlate with model size. For instance, smaller models like Llama 3 8b, as well as larger ones like Llama 3 70b and GPT-4, demonstrated a high degree of compliance with instructions to output only the answer. Conversely, other models, such as the 2 billion parameter Phi-2 and the seemingly large Mistral Large and Claude 3 Opus, frequently struggled to adhere to the desired format, often opting to provide reasons and explanations for their provided label.
To address these challenges, we developed custom parsers for each task:
@@ 2.8 Additional Considerations
In addition to the evaluation metrics and parsing strategies discussed previously, there are a few additional considerations worth noting.
First, despite our efforts in prompt engineering, certain LLM s, such as Mistral Medium and Claude 3 Opus, consistently struggled to adhere to the instructions provided in the prompts across various tasks. To mitigate this issue, we employed modified prompts in some experiments, primarily by incorporating the "repetition line" from prompt SEV-4, which explicitly reiterated the instruction to provide a concise answer without explanation. This modification generally improved the adherence of these models to the desired output format.
@@
Third, during evaluation, we compared two approaches: setting a maximum output token limit of 2 tokens versus allowing the models to respond freely without truncation. The rationale behind the former approach was that it would likely include the keyword for the diagnosis we were seeking. Conversely, for models that did not adhere well to prompt instructions, the diagnosis might appear later in their response. Our findings revealed that the longer a model’s response, the more likely it was to become invalid due to the possibility of it using ’yes’ or ’no’ when providing explanations for its decision. Consequently, we decided to impose a response length limit of 2 tokens for all models tested. Fig 2 presents a summary of each model’s invalid response percentage.
The GPT family showed strong performance, with an average invalid response rate of 0.45%. GPT-4 demonstrated impressive performance, with a reduced invalid % rate to 0.02%. However, this positive trend reversed with GPT-4-Turbo, where the rate increased to 1.64%. The subsequent GPT-4o and GPT-4o mini versions returned to low invalid rates of 0.22% and 0.0%, respectively.
As for the Llama family, Llama 3 represented a marginal improvement over Llama 2’s already impressive invalid percentage numbers (0.43% to 6.37%). However, the most recently released Llama 3.1 models perform significantly worse potentially due to their ethical guards [59]. Llama 3.1 70b is not affected as much, showing a 0.42% worse performance than Llama 3 70b, which had a 0.37% invalid rate. However, the other two models, and especially the Llama 3.1 405b model, are significantly worse, with the Llama 3.1 405b reaching a 9.0% invalid rate.
The Mistral family showed steady improvement in following instructions, starting with Mistral 7b at 12.29% and decreasing to nearly 0% for both Mistral Nemo 7b and Mistral Large 2. Despite this overall positive trend, Mistral Large unexpectedly had a high invalid response rate of 19.68%, which was later corrected in Mistral Large 2, bringing the rate back down to 0%.
The Gemma family displayed a stable pattern in invalid response rates, with most models performing consistently across iterations. The notable exception was Gemma 2b, which, likely due to its smaller size, recorded the highest invalid rate at 33.53%. Surprisingly, Gemma 7b maintained a low rate of 1.88%, similar to the larger models like Gemma 1.0 Pro (1.37%), Gemma 2 9b (0.52%), and Gemma 27b (0.35%).
The Phi, Qwen, and Claude families exhibited varied performances, with the Phi models showing a range of invalid response rates from 0.07% to 27%, with Phi-3-medium being the best. The Qwen 2 72b model had a moderate invalid response rate of 1.46%. The Claude family displayed significant variability, with rates ranging from 0.8% to 20.49%. Surprisingly Claude 3 Opus showed worse performance compared to the significantly cheaper Clause 3 Haiku, with both being significantly worse than Claude 3 Sonnet.
2022-11-282023-03-142023-07-182023-09-272023-11-142023-12-092023-12-122023-12-232024-02-212024-02-262024-03-142024-04-182024-05-132024-06-062024-06-202024-06-272024-07-022024-07-182024-07-232024-07-240.010.010.010.010.10.10.10.1111110101010100100100100Claude 3 HaikuClaude 3 SonnetClaude 3.5 SonnetClaude 3 OpusGemma 2bGemma 7bGemma 2 9bGemma 2 27bGemini 1.0 ProGPT-4o miniGPT-3.5-TurboGPT-4oGPT-4-TurboGPT-4Llama 3 8bLlama 2 13bLlama 2 70bLlama 3 70bLlama 3.1 8bLlama 3.1 70bLlama 3.1 405bMistral 7bMistral NeMoMixtral 8x7bMixtral 8x22bMistral MediumMistral LargeMistral Large 2Phi-2Phi-3-miniPhi-3-smallPhi-3-mediumQwen 2 72bRelease Dateof Invalid Responses (Log Scale)
Figure 2: Invalid Response % Statistics Sorted by model release date
@@ 3 Results & Discussion
In this section, we present the results of our experiments, analyzing the performance of various LLM s across the three tasks defined in Section 2. We begin by examining the performance variability between runs across different LLM s using consistent parameters for all experiments. This analysis establishes a foundation for assessing the reliability and reproducibility of the presented results, determining whether observed improvements between prompts are valid and significant or merely due to chance. Next, we delve into the results of the ZS binary disorder detection task (Section 3.3), followed by the ZS/FS disorder severity evaluation task (Section 3.4), and conclude with the ZS psychiatric knowledge assessment task (Section 3.5).
@@
In our evaluation of the LLM s across the three distinct tasks, we observed a wide range of performance capabilities and several noteworthy trends. In the ZS binary disorder detection task, GPT-4, GPT-4o, Mistral NeMo and Llama 2 70b consistently outperformed other models. Additionally, the choice of prompt significantly influenced model performance, with some models favoring more structured prompts while others performed better with open-ended ones. Notable mentions in this category include the Gemma 7b model which achieved 76.70% (BIN-4) accuracy on Dreaddit, which is significantly higher than the previous BIN-1 best of 74.50% of the Mistral NeMo, Mistral NeMo for competing with models much larger than it in size.
@@
The ZS/few-shot disorder severity evaluation task revealed that most models benefited from few-shot learning, where they were provided with a few examples before making predictions. However, the degree of improvement varied across models and datasets, highlighting the need for careful consideration of both model architecture and data characteristics. Finally, in the ZS psychiatric knowledge assessment task, we found that the more recent models generally performed better, compared to older, much larger models (e.g. Llama 3 70b and Phi-3-medium competing with GPT-4, GPT-4-Turbo, Claude 3.5 Sonnet, and others). However, with models released around the same time period, model size seemed to mostly be the deciding factor (Llama 3.1 models are excluded from this finding due to their tendency to filter/not respond to our prompts due to their sensitive mental health nature).
3.1 Performance Variability Experiment
In the performance variability experiment, we evaluate the inherent performance variability of each LLM on the same task. The experiment involved repeatedly evaluating each model on the same dataset five times with the standard parameters we typically use (e.g., temperature 0). This approach allows us to approximate the natural performance fluctuations, enabling us to judge whether any observed performance enhancement brought about by changes to the prompt or other factors is a meaningful improvement or a random fluctuation due to the model’s inherent variability. The experiment was conducted on 1000 fairly sampled instances from the DEPTWEET dataset.
3.2 Performance Variability Experiment
In the performance variability experiment, we evaluate the inherent performance variability of each LLM on the same task. The experiment involved repeatedly evaluating each model on the same dataset five times with the standard parameters we typically use (e.g., temperature 0). This approach allows us to approximate the natural performance fluctuations, enabling us to judge whether any observed performance enhancement brought about by changes to the prompt or other factors is a meaningful improvement or a random fluctuation due to the model’s inherent variability. The experiment was conducted on 1000 fairly sampled instances from the DEPTWEET dataset.
@@ 3.2 Performance Variability Experiment
The results in Table 7 show that most models exhibited performance fluctuations of less than 0.5%, with the exception of Mixtral 8x22b, which showed a 1.6% deviation. Models not listed in the table had a standard deviation of 0%, indicating no performance variability across the five runs. These results suggest that observed changes in model performance (as long as they are larger than the observed fluctuations) are likely due to our modifications (e.g., to the prompt utilized) rather than inherent model variability.

### v2

@@ 3.4 Models
Our model selection encompassed both closed-source models, accessible through APIs, and open-source models, which offer greater transparency and flexibility for research purposes. We included models from various families, including OpenAI’s GPT series, Anthropic’s Claude models, Mistral AI’s models, Google’s models, Meta’s Llama series, and finally, Phi-2 and MentaLlama, a model specifically fine-tuned on mental health-related data. Additionally, we attempted to include Gemini 1.5 Pro in our evaluation, however, due to rate limit constraints and filtering mechanisms challenges mentioned in Section 7.1, we were not able to include it.
Table 3 provides a summary of the models used in our experiments, along with their respective sizes, the APIs through which we accessed them, and the costs for running our experiments.
Table 3: Summary of large language models evaluated in this analysis and cost of running our experiments.
Model Size (Parameters) Source APIService Provider Experiment Costs (USD)
Claude 3 Haiku [anthropic2024claude] Proprietary Anthropic Claude API \$2.3
Claude 3 Sonnet [anthropic2024claude] Proprietary Anthropic Claude API \$9.8
Claude 3 Opus [anthropic2024claude] Proprietary Anthropic Claude API \$33.8
Claude 3.5 Sonnet [anthropic2024claude35] Proprietary Anthropic Claude API \$6.8
Gemma 2b [team2024gemma] 2B Google Nvidia NIM \$0.8
Gemma 7b [team2024gemma] 7B Google Groq \$0.0
Gemma 2 9b [team2024gemma2] 9B Google Nvidia NIM \$1.5
Gemma 2 27b [team2024gemma2] 27B Google Nvidia NIM \$3.8
Gemini 1.0 Pro [team2023gemini] Proprietary Google Gemini API \$0.0
GPT-4o mini [openai2024gpt4omini] Proprietary OpenAI OpenAI API \$0.8
GPT-3.5-Turbo [openai2023gpt35turbo] 175B OpenAI OpenAI API \$3.8
GPT-4o [openai2024gpt4o] Proprietary OpenAI OpenAI API \$11.3
GPT-4-Turbo [achiam2023gpt] Proprietary OpenAI OpenAI API \$22.5
GPT-4 [achiam2023gpt] Proprietary OpenAI OpenAI API \$65.3
Llama 3 8b [dubey2024llama] 8B Meta Groq \$0.0
Llama 3.1 8b [dubey2024llama] 8B Meta Together AI \$1.5
Llama 2 13b [touvron2023llama2] 13B Meta Together AI \$1.5
Llama 2 70b [touvron2023llama2] 70B Meta Together AI \$6.0
Llama 3 70b [dubey2024llama] 70B Meta Groq \$0.0
Llama 3.1 70b [dubey2024llama] 70B Meta Together AI \$4.5
Llama 3.1 405b [dubey2024llama] 405B Meta Together AI \$25.5
Mistral 7b [jiang2023mistral] 7B Mistral AI Together AI \$1.5
Mistral NeMo [mistral2024nemo] 12B Mistral AI Mistral AI API \$1.5
Mixtral 8x7b [jiang2024mixtral] 45B (13B) Mistral AI Groq \$0.0
Mixtral 8x22b [jiang2024mixtral] 141B (39B) Mistral AI Together AI \$6.8
Mistral Medium [mistral2023la] Proprietary Mistral AI Mistral AI API \$17.3
Mistral Large [mistral2024large] Proprietary Mistral AI Mistral AI API \$18.0
Mistral Large 2 [mistral2024large2] Proprietary Mistral AI Mistral AI API \$15.0
Phi-2 [javaheripi2023phi2] 2.7b Phi-1 Labs Nvidia NIM \$0.8
Phi-3-mini [abdin2024phi] 3.8B Phi-1 Labs Nvidia NIM \$0.8
Phi-3-small [abdin2024phi] 7B Phi-1 Labs Nvidia NIM \$1.5
Phi-3-medium [abdin2024phi] 14B Phi-1 Labs Nvidia NIM \$2.3
Qwen 2 72b [yang2024qwen2] 72B Alibaba Cloud Together AI \$4.5
Note 1: The Mixtral 8x7b and 8x22b models have two sizes because the total size and the number of parameters utilized at any point in inference are different.
Note 2: Gemini 1.0 Pro API and Groq models were available for free at the time these experiments were conducted. These are approximate costs, accounting only for the final runs presented in this paper and do not include any experimentation costs, which would drastically increase the overall expenditure.
@@ 3.5 Prompt Templates
The prompt, or instruction, given to an LLMplays a significant role in its performance on a given task. Prior research has shown that even slight variations in prompt wording can significantly impact the quality of LLMoutputs [salinas2024butterfly]. Therefore, we dedicated considerable effort to crafting effective prompts tailored to each task in our study.
To optimize LLMperformance on social media post analysis, we employed prompts that encouraged the models to adopt the role of a psychiatrist. Role-playing has proven effective in eliciting relevant and contextually appropriate responses from LLMs [kong2023better, zheng2023helpful]. We also utilized the concept of prompt templates, which consist of a fixed structure with variable elements that can be adapted for different tasks. This approach ensured consistency across experiments while minimizing prompt-based bias.
@@ 3.5.1 Binary Disorder Detection Prompt Templates
As shown in Table 4, Prompts BIN-1 and BIN-2 utilized a straightforward structure, first instructing the LLMto assume the role of a psychiatrist and then directly asking whether the poster of a given social media post exhibited a specific disorder. We included a final instruction to elicit a "yes" or "no" response for easier parsing of the LLMoutput.
@@
[Task]
Act as a psychologist. Label the writer of the provided social media post as ’depressed’
or ’not depressed’ according to provided guidelines.
[Guidelines]
- Concise Response: Respond only with ’Yes’ (depressed) or ’No’ (not depressed).
- No Explanations: Don’t provide explanations for your assessment.
- Ambiguity: If the post is unclear, choose the most probable label.
[Post]
@@
Prompts BIN-3 and BIN-4 introduced a more structured format, explicitly separating the prompt into "Task," "Guidelines," and "Post" sections. The guidelines emphasized the desired output format (yes/no), discouraged explanations, and instructed the LLMto choose the most probable label in cases of ambiguity. This structure aimed to improve the LLMs’ adherence to the task instructions and reduce the incidence of irrelevant or unparseable responses. The inspiration for this structure was from Gemini, as this is the structure used for its training [team2023gemini].
@@ 3.5.2 Disorder Severity Evaluation Prompt Templates
Figure 1: Disorder Severity Evaluation Task Evolution
@@ 3.6 Evaluation Metrics
For the disorder severity evaluation task, we employed two metrics to assess the performance of LLMs: BAand MAE. BA, in this context, refers to the proportion of posts for which the LLMpredicted the exact severity level assigned by human annotators. It serves as a measure of how often the model gets the severity classification exactly right. On the other hand, MAEquantifies the average magnitude of the errors in the predicted severity ratings. It provides a complementary perspective on model performance by assessing how close, on average, the predicted severity levels are to the true levels, even if they are not exactly correct. A model could potentially improve in one metric while worsening in the other.
@@ 3.7 Parsing Model Outputs
To manage this output variability and facilitate the extraction of relevant information, we developed custom parsers designed for each task.
@@ 3.8 Additional Considerations
In addition to the evaluation metrics and parsing strategies discussed previously, there are a few additional considerations worth noting. First, despite our efforts in prompt engineering, certain LLMs, such as Mistral Medium and Claude 3 Opus, struggled to adhere to instructions across tasks. During the design of the Severity prompts (Section 3.5.2), we observed that including a repetition line (“(Please return only the label (Yes or No) and no other text)”) improved compliance. To address similar issues in the binary classification task, we created a modified prompt, BIN1.1, by appending this repetition line to the original binary prompt (see Table 12). This adjustment improved adherence to output formatting requirements in subsequent experiments. We also conducted a small investigation into the effect of this repetition line on model performance (see Section 6.1).
@@
Third, during evaluation, we compared two approaches: setting a maximum output token limit of 2 tokens versus allowing the models to respond freely without truncation. The rationale behind the former approach was that it would likely include the keyword for the diagnosis we were seeking. Conversely, for models that did not adhere well to prompt instructions, the diagnosis might appear later in their response. Our findings revealed that the longer a model’s response, the more likely it was to become invalid due to the possibility of it using ’yes’ or ’no’ when providing explanations for its decision. Consequently, we decided to impose a response length limit of 2 tokens for all models tested. A detailed analysis of the invalid response rates across different models, which informed this decision, is presented in Section 4.2.3.
@@ 4 Results & Discussion
In this section, we present the results of our experiments, analyzing the performance of various LLMs across the three tasks defined in Section 3. We first present the results of our performance variability assessment (Section 4.1), which provides insights into the inherent reliability of each model and helps contextualize the significance of performance differences observed in subsequent tasks. Next, we delve into the results of the ZSbinary disorder detection task (Section 4.2), followed by the ZS/FSdisorder severity evaluation task (Section 4.3), and the ZSpsychiatric knowledge assessment task (Section 4.4). We then present an initial exploration of fine-tuning for the severity evaluation task (Section 4.5). Finally, we conclude with a brief discussion of the study’s limitations (Section 4.6), which will be explored in greater detail in Section 7.1.
@@
In our evaluation of the LLMs across the three distinct tasks, we observed a wide range of performance capabilities and several noteworthy trends. In the ZSbinary disorder detection task, GPT-4, GPT-4o, Mistral NeMo, and Llama 2 70b consistently outperformed other models. Additionally, the choice of prompt significantly influenced model performance, with some models favoring more structured prompts while others performed better with open-ended ones. Notable mentions in this category include the Gemma 7b model which achieved 76.70% (BIN-4) accuracy on Dreaddit, which is significantly higher than the previous BIN-1 best of 74.50% achieved by Mistral NeMo. Mistral NeMo is also notable for competing with models much larger than its size.
@@
The ZS/FSdisorder severity evaluation task revealed that most models benefited from FSlearning, where they were provided with a few examples before making predictions. However, the degree of improvement varied across models and datasets, highlighting the need for careful consideration of both model architecture and data characteristics. Finally, in the ZSpsychiatric knowledge assessment task, we found that the more recent models generally performed better, compared to older, much larger models (e.g. Llama 3 70b and Phi-3-medium competing with GPT-4, GPT-4-Turbo, Claude 3.5 Sonnet, and others). However, with models released around the same time period, model size seemed to mostly be the deciding factor (Llama 3.1 models are excluded from this finding due to their tendency to filter/not respond to our prompts due to their sensitive mental health nature).
4.1 Performance Variability Results
Table 8 presents the results of our performance variability assessment, as described in Section 3.1. The table shows the standard deviation of accuracy for each model across five repeated evaluations on the DEPTWEET dataset.
The results indicate that most models exhibited low performance variability, with standard deviations below 0.5%. Specifically, models such as Gemma 2 9b, Gemma 2 27b, GPT-3.5-Turbo, Llama 2 70b, Llama 3 70b, Phi-3-small, Phi-3-medium, and Qwen 2 72b fall within this range. Mixtral 8x22b showed a higher variability with a standard deviation of 1.6%. Models unlisted in the table demonstrated 0% standard deviation.
These results are crucial for interpreting the significance of performance differences observed in the subsequent task results. The consistently low variability of most models suggests that performance changes exceeding approximately 0.5% (or 1.6% for Mixtral 8x22b) are likely due to experimental manipulations (e.g., prompt variations) and not random model fluctuations. This allows us to more confidently analyze the impact of different factors on model performance in the tasks that follow.